OpenStudy (anonymous):

solve the quadratic equation using the quadratic formula x^2+7x+3=0

OpenStudy (anonymous):

do you know how to factor?

OpenStudy (anonymous):

yeah kinda

OpenStudy (mertsj):

Here is the quadratic formula: \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

yeah i know it thanks lol i already plugged them in

OpenStudy (mertsj):

In your problem a = 1, b = 7, c = 3 Plug in those values and simplify.

OpenStudy (anonymous):

i got 

OpenStudy (anonymous):

yeah how did you get 37?

OpenStudy (mertsj):

You should get: \[x=\frac{-7}{2}\pm\frac{\sqrt{37}}{2}\]

OpenStudy (mertsj):

b^2=49 4(1)3)=12 49-12=37

OpenStudy (anonymous):

Ohh, Okay thank you, but can you please show me how you got 37

OpenStudy (mertsj):

\[\sqrt{7^2-4(1)(3)}=\sqrt{49-12}=\sqrt{37}\]

OpenStudy (anonymous):

okay thank :) i think i get it but is that simplied completley

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

so those are the soulutions

OpenStudy (mertsj):

yes